The Singer Division of Patio Enterprises currently earns $2.87 million and has divisional assets of $20.5 million. The division manager is considering the acquisition of a new asset that will add to profit. The investment has a cost of $3,447,000 and will have a yearly cash flow of $858,000. The asset will be depreciated using the straight-line method over a six-year life and is expected to have no salvage value. Divisional performance is measured using ROI with beginning-of-year net book values in the denominator. The company’s cost of capital is 12 percent. Ignore taxes.

Required:

a. What is the divisional ROI before acquisition of the new asset? (Enter your answer as a percentage rounded to 1 decimal place (i.e., 32.1).)

b. What is the divisional ROI in the first year after acquisition of the new asset? (Enter your answer as a percentage rounded to 1 decimal place (i.e., 32.1).)
Answer #1

a. ROI = Operating profit / Net book value

= 2.87 / 20.5 = 14%

b. Depreciation = 3447000 /6=574500

Profit on new asset = 858000-574500=283500

ROI = (2870000+283500) /(20500000+3447000) = 13.2%